OpenStudy (anonymous):

Find the derivative of f(x) = 7x + 9 at x = 6.

hartnn (hartnn):

do you know how to find derivatives ? derivative of x^n is ?

OpenStudy (anonymous):

I do not remember @hartnn

OpenStudy (precal):

use the power rule

OpenStudy (anonymous):

Which is

hartnn (hartnn):

\(\Large \dfrac{d}{dx} x^n = n x^{n-1}\)

hartnn (hartnn):

for\( 7x =7x^1 \) the exponent is 1, so plug in n=1

OpenStudy (anonymous):

so that would be 7

hartnn (hartnn):

yes, and derivative of constant = 0 so derivative of 7x + 9 is simply 7

OpenStudy (anonymous):

Okay so for 4/x and x=2, would it be 4?

hartnn (hartnn):

nopes, 4/x is \(\Large 4 x^{-1}\) so you will have to plug in x= -1 in that formula

OpenStudy (anonymous):

-4?

hartnn (hartnn):

no, what did you get after plugging in n =-1 in that formula ? \(\Large (-1)x^{-1-1} = ... ?\)

OpenStudy (anonymous):

wait plug in -1 for x in 4/x? 4/-1

hartnn (hartnn):

no, plug in -1 for n in the formula for derivative

OpenStudy (anonymous):

Haha I'm confused now, sorry

hartnn (hartnn):

you want derivative of \(x^{-1}\) and you have the formula as \(\Large \dfrac{d}{dx} x^n = n x^{n-1}\)

hartnn (hartnn):

so naturally, you will select n as -1

hartnn (hartnn):

\(\Large \dfrac{d}{dx} x^{-1} = (-1)x^{-1-1} =...?\)

OpenStudy (anonymous):

0?

hartnn (hartnn):

no.. what is -1-1 = .. ?

OpenStudy (anonymous):

-2 lol sorry

hartnn (hartnn):

yeah, so derivative of x^-1 = -1 x^-2 derivative of 4 x^-1 will be -4 x^-2 = -4/x^2 now you can plug in x=2 here

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

wait i read that wrong

OpenStudy (anonymous):

-1

hartnn (hartnn):

-4/2^2 = -4/4 = -1 yes! you're correct :)

OpenStudy (anonymous):

Thank you (:

hartnn (hartnn):

welcome ^_^